# silent-chen.github.io/DGE

### v1

@@ 5.3 Partial Editing
Finally, we showcase some local editing results on the Horns scene from LLFF [68] dataset. Firstly, we segment the Gaussians and get the mask of the Gaussians. Then, we render images from different views and edit those views with our proposed DGE. The results are realistic and only applied to the segmented skeleton, demonstrating our method’s ability of partial editing.
@@ Appendix 0.A Overview
  • Additional Editing Results (Appendix 0.B). We provide additional qualitative examples of our method, including more editing prompts and partial editing results on complex scenes with many objects.
  • Additional Comparisons (Appendix 0.C). We provide in-depth comparisons of our approach and other update schemes and further comparisons with prior work, GaussianEditor and ViCA-NeRF.
  • Additional Details (Appendix 0.D). We provide more details regarding the experimental implementation and evaluation protocol.
  • Social Impact and Limitations (Appendix 0.E). We discuss the social impact and limitations of our work.
Appendix 0.B Additional Editing Results
@@
Results with Additional Prompts.
@@
Results with Partial Editing.
@@
Appendix 0.C Additional Comparisons
Multi-view Consistent Editing vs. Iterative Dataset Updates.
@@ Appendix 0.C Additional Comparisons
Extended Comparisons with GaussianEditor.
@@
Figure 12: Qualitative Comparison with ViCA-NeRF. We use the official implementation from its GitHub repository. ViCA-NeRF results in blurry editing results when asked to turn the man into “mosaic sculpture” or “Iron Man”. It also fails when given prompts such as “Turn him into a Spider Man with a mask” or “Make him wear a Venetian mask”.
Method 3D CLIP CLIP Directional Avg. Editing
Model Similarity Similarity Time
ViCA-NeRF  [28] NeRF 0.204 0.44 $\sim$ 28min
Ours GS 0.222 0.66 $\sim$ 4min
Table 3: Comparison with ViCA-NeRf. Our DGE achieves better performance with much less time.
Comparison with ViCA-NeRF.
We provide an additional comparison between our work and ViCA-NeRF [19], which also focuses on editing different views consistently. We use the official implementation provided in their GitHub repository. We highlight some key differences as below. ViCA-NeRF uses 3D model depth information to project features from key views into others using a blending module. Our approach, however, takes inspiration from [24], which injects attention values from keyframes into other frames. In addition, we incorporate further 3D priors to constrain this process with epipolar lines, minimizing the corresponding matching errors. This leads to clearer results compared to ViCA-NeRF’s, which can sometimes produce blurry edits or fail. We have conducted both qualitative and quantitative evaluations, showing the superiority of our method in achieving more accurate and higher-quality edits in only a fraction of the time. Examples and detailed comparisons are in Figure 12 and Table 3. As shown in Figure 12, ViCA-NeRF results in blurry editing results and fails to produce meaningful editing given prompts such as “Turn him into spider man with a mask”. The quantitative results in Table 3 also suggest that our method surpasses ViCA with less time and higher quality.
Appendix 0.D Additional Details
0.D.1 Implementation Details
@@ 0.D.2 Evaluation Prompts
Face “A man with curly hair in a grey jacket.” “A man with curly hair in a grey jacket with a Venetian mask.” “Give him a Venetian mask.”
Face “A man with curly hair in a grey jacket.” “A man with curly hair in a checkered cloth.” “Give him a checkered jacket.”
Face “A man with curly hair in a grey jacket.” “A spider man with a mask and curly hair.” “Turn him into spider man with a mask.”
Bear “A stone bear in a garden.” “A robotic bear in the garden.” “Make the bear look like a robot.”
Bear “A stone bear in a garden.” “A panda in the garden.” “Make the bear look like a panda.”
Bear “A stone bear in a garden.” “A bear with rainbow color” “Make the color of the bear look like rainbow.”
Person “A man standing next to a wall wearing a blue T-shirt and brown pants.” “A man looks like a mosaic sculpture standing next to a wall.” “Make the man look like a mosaic sculpture.”
Person “A man standing next to a wall wearing a blue T-shirt and brown pants.” “A man wearing a shirt with a pineapple pattern.” “Make the person wear a shirt with a pineapple pattern.”
Person “A man standing next to a wall wearing a blue T-shirt and brown pants.” “An Iron Man stands to a wall.” “Turn him into Iron Man.”
Person “A man standing next to a wall wearing a blue T-shirt and brown pants.” “A robot stands to a wall.” “Turn the man into a robot.”
@@
Appendix 0.E Social Impact and Limitations

### v2

@@ 5.3 Partial Editing
Finally, in Figure 6, we showcase partial (local) editing results on the Horns scene from the LLFF [60] dataset. First, we segment 2D renders to get 2D masks of the skeleton we want to edit. Then, we unproject the 2D masks to 3D getting the mask of the Gaussians. As last, we render images from different views and edit those views with our proposed DGE. The results are realistic and only applied to the segmented skeleton, demonstrating our method’s ability of partial editing. We provide additional partial editing results in the supplementary material.
@@ 6 Conclusion
  • Additional Editing Results (Appendix 0.A). We provide additional qualitative examples of our method, including more editing prompts and partial editing results on complex scenes with many objects.
  • Additional Comparisons (Appendix 0.B). We provide in-depth comparisons of our approach and other update schemes and further comparisons with prior work, GaussianEditor.
  • Additional Details (Appendix 0.C). We provide more details regarding the experimental implementation and evaluation protocol.
  • Social Impact and Limitations (Appendix 0.D). We discuss the social impact and limitations of our work.
Appendix 0.A Additional Editing Results
@@
0.A.0.1 Results with Additional Prompts.
@@
0.A.0.2 Results with Partial Editing.
@@
Appendix 0.B Additional Comparisons
0.B.0.1 Multi-view Consistent Editing vs. Iterative Dataset Updates.
@@ Appendix 0.B Additional Comparisons
0.B.0.2 Extended Comparisons with GaussianEditor.
@@
Appendix 0.C Additional Details
0.C.1 Implementation Details
@@ 0.C.2 Evaluation Prompts
Face “A man with curly hair in a grey jacket” “A man with curly hair in a grey jacket with a Venetian mask” “Give him a Venetian mask”
Face “A man with curly hair in a grey jacket” “A man with curly hair in a checkered cloth” “Give him a checkered jacket”
Face “A man with curly hair in a grey jacket” “A spider man with a mask and curly hair” “Turn him into spider man with a mask”
Bear “A stone bear in a garden” “A robotic bear in the garden” “Make the bear look like a robot”
Bear “A stone bear in a garden” “A panda in the garden” “Make the bear look like a panda”
Bear “A stone bear in a garden” “A bear with rainbow color” “Make the color of the bear look like rainbow”
Person “A man standing next to a wall wearing a blue T-shirt and brown pants” “A man looks like a mosaic sculpture standing next to a wall” “Make the man look like a mosaic sculpture”
Person “A man standing next to a wall wearing a blue T-shirt and brown pants” “A man wearing a shirt with a pineapple pattern” “Make the person wear a shirt with a pineapple pattern”
Person “A man standing next to a wall wearing a blue T-shirt and brown pants” “An Iron Man stands to a wall” “Turn him into Iron Man”
Person “A man standing next to a wall wearing a blue T-shirt and brown pants” “A robot stands to a wall” “Turn the man into a robot”
@@
Appendix 0.D Social Impact and Limitations